OpenStudy (anonymous):

The matrix C=[ ] was used to encode a phrase to [ ]. Find C^-1 and use it to decode the matrix. I will post a pic of the matrices in the comments. Please help!

OpenStudy (anonymous):

@jim_thompson5910 Could you help with this last one as well?

jimthompson5910 (jim_thompson5910):

do you know how to find the inverse of a 2x2 matrix?

OpenStudy (anonymous):

I'm not exactly sure

OpenStudy (anonymous):

Multiply it by its negative?

jimthompson5910 (jim_thompson5910):

one sec

jimthompson5910 (jim_thompson5910):

If \[\Large X = \begin{bmatrix}a & b\\ c & d \end{bmatrix}\] Then \[\Large X^{-1} = \frac{1}{ad-bc}\begin{bmatrix}d & -b\\ -c & a \end{bmatrix}\]

jimthompson5910 (jim_thompson5910):

Does that formula look familiar?

OpenStudy (anonymous):

Yes it does.

jimthompson5910 (jim_thompson5910):

So what is the inverse of the given 2x2 matrix?

OpenStudy (anonymous):